OpenStudy (anonymous):

What is the difference between y=log (100x) and y= 2+logx Graphically they are the identical, but I do not know algebraically how there are differently or identical. Need help plz.

OpenStudy (anonymous):

ok its the same thing

OpenStudy (anonymous):

and I will show u why algebraically

OpenStudy (anonymous):

Are u still there?

OpenStudy (anonymous):

yea im here

OpenStudy (anonymous):

Ohh sorryyy. I was not getting messages today There is a property of logs: log(a*b)=log(a)+log(b)

OpenStudy (anonymous):

y=log(100x) Using the property of logs we have y=log(100)+log(x) so y =2+log(x)

OpenStudy (anonymous):

If you were not aware log(100)=2 umm may confuse you but if you know logs than its simple 10^2=100